# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v2

@@ Abstract
While many advances in time series models focus exclusively on numerical data, research on multimodal time series, particularly those involving contextual textual information commonly encountered in real-world scenarios, remains in its infancy. With recent progress in large language models and time series learning, we revisit the integration of paired texts with time series through the Platonic Representation Hypothesis [15], which posits that representations of different modalities converge to shared spaces. In this context, we identify that time-series-paired texts may naturally exhibit periodic properties that closely mirror those of the original time series. Building on this insight, we propose a novel framework, Texts as Time Series (TaTS), which considers the time-series-paired texts to be auxiliary variables of the time series. TaTS can be plugged into any existing numerical-only time series models and enable them to handle time series data with paired texts effectively. Through extensive experiments on both multimodal time series forecasting and imputation tasks across benchmark datasets with various existing time series models, we demonstrate that TaTS can enhance predictive performance without modifying model architectures. Code available at https://github.com/iDEA-iSAIL-Lab-UIUC/TaTS.
@@ 1 Introduction
Figure 1: Mean Square Error of modeling frameworks of time series with paired texts on the Environment dataset.
Time series modeling plays an important role in a wide range of real-world applications, including domains of finance [46, 17, 74], healthcare [65, 13], climate [10], and energy systems [24, 27, 18, 20]. While extensive research has focused on approaches that rely solely on the numerical values of time series [42, 33, 75, 37, 58, 53, 26], real-world scenarios often involve additional modalities that co-occur with the time series and can provide valuable complementary information [6, 44, 25, 54, 19, 72, 71].
In scenarios like pandemic policymaking, economic planning, or investment strategies, textual information can provide explanations, updates, or external factors that influence the underlying numerical patterns. However, research on effectively leveraging data from other modalities paired with time series remains in its early stages. In this work, we focus on time-series paired with texts at each timestamp, a common data format in real-world applications where textual descriptions are associated with time series at each timestamp in a parallel manner, as illustrated in Figure 3 (left). For instance, during a pandemic, infection rates are often accompanied by textual information like government announcements and news reports in real-time [5]. On the one hand, numerical-only models overlooks valuable contextual information that may influence or explain the patterns in the time series. On the other hand, the current state-of-the-art approach [35] treats the textual modality as “bag-of-texts", disregarding the unique positional characteristics that time-series-paired texts may inherently possess. Such limitations raise a pivotal question:
@@
In this paper, motivated by the Platonic Representation Hypothesis (PRH) [15], we pioneer the exploration of effectively leveraging paired texts to enrich time series analysis. We identify an intriguing phenomenon, which we term “Chronological Textual Resonance” (CTR): depending on data quality, time-series-paired texts may exhibit periodic patterns that closely reflect the temporal dynamics of their corresponding numerical time series. Notably, despite variations in words and expressions, the hidden representations of two texts associated with time series points separated by a period may demonstrate high similarities, revealing a deeper alignment between textual and numerical modalities. We attribute this phenomenon to the fact that the paired texts are highly sensitive to and inherently evolve in response to the dynamics of the time series itself. Further, we introduce TT-Wasserstein, a new metric designed to measure the level of CTR and quantify alignment quality.
Building on these insights, we propose Texts as Time Series (TaTS), a simple yet effective framework for integrating paired texts to enhance multimodal time series modeling. Previous studies have shown that different variables in one multivariate time series exhibit similar periodicity properties [67, 54, 62], and CTR suggests that time-series-paired texts follow a similar pattern. This observation implies that paired texts can be considered as special auxiliary variables to augment the original time series. Motivated by this, TaTS first transforms the paired textual information into a lower-dimensional representation, then combines the original time series with the textual representations as new variables to form an augmented time series. This augmented time series is subsequently fed into existing time series models, allowing them to capture both numerical and textual dynamics. TaTS offers two key benefits: (i) it effectively captures the evolving positional characteristics of texts paired with a time series by incorporating their projected representations into time series models; and (ii) it functions as a plug-in module, maintaining compatibility with existing time series models. Empirically, through extensive experiments over various benchmark datasets and multiple existing time series models, the proposed TaTS achieves state-of-the-art performance on both forecasting and imputation tasks. Notably, we observe that a higher CTR level (i.e., a lower TT-Wasserstein) correlates with greater improvements compared to numerical-only modeling. In summary, our contributions are:
@@
  • Experiments on diverse benchmark datasets and multiple existing time series models demonstrate TaTS’s superior performance without requiring modifications to model architectures.
@@ 2 Preliminary
where $\mathbf{e}_{t}$ is the encoded representation of the text at time $t$, $d_{\text{text}}$ is the dimensionality of the textual embedding, and $\theta_{\text{text}}$ are the parameters of the encoder. In this work, we leverage pre-trained large language models to encode the texts by applying pooling over embeddings of individual tokens.
@@ 3 Chronological Textual Resonance (CTR)
The Platonic Representation Hypothesis. PRH [15] posits that different modalities describing the same object converge towards a shared, latent representation. Extending this hypothesis, if time series and paired text both describe the same changing event, their representations can be seen as dynamic projections from a common underlying source, and should exhibit similar periodic properties.
To illustrate this hypothesis for time series with paired texts, we analyze three real-world datasets [35], including (i) Economy: The time series represents trade data of the U.S., while the texts describe the general economic conditions of the country. (ii) Social Good: The time series captures the unemployment rate in the U.S., and the texts include detailed unemployment reports. (iii) Traffic: The time series reflects monthly travel volume trends from the U.S. Department of Transportation, with corresponding texts derived from traffic volume reports issued by the same department.
For each dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, we employ the Fourier Transform [40, 47] to analyze the frequency components of time series data, allowing us to identify its dominant periodic components, as illustrated by the blue curves in Figure 2. Furthermore, to examine the periodicity of texts, we embed each $s_{t}\in\mathbf{S}$ to obtain the text embedding $e_{t}$ at timestamp $t$. Then we compute their lag-similarity, defined as $d_{l}=\sum_{t}\mathrm{cos}(e_{t},e_{t+L})$ where $L$ is the lag and $\mathrm{cos}(\cdot,\cdot)$ represents the cosine similarity. If the text embeddings exhibit a significant periodic pattern, the lag-similarity $d_{l}$ will also fluctuate periodically as the lag $l$ increases (proof in Proposition A.1). Finally, we identify the major frequencies (those with the largest amplitudes) of the texts by applying FFT to the text lag-similarity, and mark them with red dashed lines, as shown in Figure 2. Detailed process is provided in Appendix A. We find that, for these datasets, the major frequencies of the paired texts closely match those of the time series. Specifically, the paired texts also show periodicity of 12 (frequency 0.083) for monthly sampled time series, indicating that the paired texts exhibit periodicity that is strongly aligned with the temporal dynamics of the time series.
@@
Why may CTR happen? We present three key reasons for the observed alignment in periodicity between time series and their paired texts: (i) Shared External Drivers: Both the time series and their paired texts are often influenced by common external factors, such as seasonal changes, recurring events, or societal and economic cycles. These shared drivers naturally induce periodicity in both the numerical time series and the accompanying texts. (ii) Influence of Time Series on Texts: Paired texts often serve as contextual reflections of the underlying time series, adapting and evolving in response to numerical trends. For instance, news articles or government reports accompanying economic indicators are frequently updated in response to the numerical trends, leading to a periodic alignment between the two modalities. (iii) Texts Contain Additional Variables with Aligned Periodicity: Paired texts often include additional variables that are closely related to the original time series. For example, if the time series represents economic data such as GDP, the accompanying texts may reference related variables like stock market indices or inflation rates. These related variables often exhibit periodicity patterns aligned with the time series and affect the periodicity of the paired texts, but are not included in the time series data.
Quantifying CTR level. Of course we cannot expect every time series with paired texts at each timestamp to exhibit periodicity similarity. In some cases, texts aligned with timestamps may not provide meaningful or complementary information, for example, daily lottery winning numbers. To this end, we propose a new metric for CTR, TT-Wasserstein, defined as the Wasserstein distance between the normalized spectra of time series and texts. Formally, based on Wasserstein distance $W(P,Q)$ [21], for time series with paired texts dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, after computing normalized frequencies and amplitudes $\tilde{f}_{\text{texts}},\tilde{f}_{\text{ts}},\tilde{a}_{\text{texts}},\tilde{a}_{\text{ts}}$,
[…]
Table 1: TT-Wasserstein measure of Time-MMD [35] datasets and their random shuffle.
@@ 4 Texts as Time Series
This section introduces our proposed framework, Texts as Time Series (TaTS). TaTS integrates concurrent texts by transforming them into auxiliary variables and seamlessly plugging the text-augmented time series into any existing time series model. Our approach effectively captures the temporal dynamics and interconnections between the time series and concurrent texts by unifying them within a single time series model. An overview of the proposed TaTS is illustrated in Figure 3.
4.1 Concurrent Texts are Secretly Auxiliary Variables
Previously, we elucidated three key reasons behind the CTR phenomenon, which serve as the core motivation for TaTS: mapping concurrent texts to auxiliary variables to augment the original time series. The properties of concurrent texts closely align with those of variables in a multivariate time series: similar to numerical variables, concurrent texts are influenced by shared external drivers and interact dynamically with the time series. Furthermore, mapping concurrent texts to structured variables enables capturing hidden variables embedded within the concurrent texts.
4.2 Framework and Training Objectives
Our proposed TaTS framework builds on the insight that concurrent texts behave similarly to additional variables and, therefore, can be treated in a comparable manner.
Transforming Concurrent Texts into Variables. Given the dataset $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}\}$, we first embed the texts using a text encoder $\mathcal{H}_{\text{text}}$ to obtain text embeddings $\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\in\mathbb{R}^{d_{\text{text}\times T}}$. Since the text embedding dimension $d_{\text{text}}$ is typically much larger than the number of variables in the time series, we reduce the dimensionality of the text embeddings by applying a Multi-Layer Perceptron (MLP), mapping them into a lower-dimensional space of reduced dimensionality $d_{\text{mapped}}$.
$\mathbf{z}_{t}=\text{MLP}(\mathbf{e}_{t};\theta_{\text{MLP}})\in\mathbb{R}^{d_{\text{mapped}}},$ (5)
@@ 4.2 Framework and Training Objectives
$\mathbf{U}=[\mathbf{X};\mathbf{Z}^{\intercal}]_{\text{dim=1}}\in\mathbb{R}^{T\times(N+d_{\text{mapped}})},$ (6)
@@
$\widehat{\mathbf{X}}_{T+1:T+H}=\mathcal{F}\left(\mathbf{U}_{1:T};\theta_{\text{forecast}}\right)[:N]\in\mathbb{R}^{H\times N},$ (7)
[…]
@@
$\mathcal{L}_{\text{forecast}}(\mathbf{X},\widehat{\mathbf{X}})=\frac{1}{H\cdot N}\sum_{t=T+1}^{T+H}\sum_{i=1}^{N}\left(\mathbf{X}_{t,i}-\widehat{\mathbf{X}}_{t,i}\right)^{2},$ (8)
@@ 5 Experiment
Table 2: Performance of time-series forecasting task. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework consistently and significantly enhances existing time series models to effectively handle time series with paired texts. The results are averaged across all prediction lengths. Full results are reported in Appendix E.3. Promotion (positive or negative) denotes the percentage reduction (or increase) in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
6.0% 8.8% 3.4% 9.1% 10.6% 14.4% 10.0% 9.1% 10.4% 10.9% 15.0% 17.3% 16.5% 15.8% 4.7% 6.0% 4.1% 7.1%
In this section, we empirically validate the effectiveness of the proposed TaTS framework through extensive experiments conducted on multiple datasets using a variety of existing time series models. In particular, we use GPT2 [43] encoder to embed the paired texts. We also validate the performance of TaTS with other text encoders across different datasets in section 5.2 and Appendix E.7. Implementation details are provided in Appendix D.4.
Datasets. We evaluate our framework on 9 real-world datasets [35], spanning diverse domains such as environmental sciences, social systems, and infrastructure. The datasets have sample frequencies ranging from daily to weekly and monthly. Further details about the datasets are in Appendix D.1.
Time Series Models and Baselines. To demonstrate the compatibility of TaTS with existing time series models, we integrate TaTS with 9 widely used models across different categories, including (i) Transformer-based models: iTransformer [37], PatchTST [39], Crossformer [67], Autoformer [56], Informer [75] and Transformer [50]. (ii) Linear models: DLinear [63]. (iii) Frequency-based models: FEDformer [77], FiLM [76]. We also compare our TaTS framework with (a) traditional covariate-based methods: N-BEATS [41], N-HiTS [4]; (b) convolution-based method TCN [1]; (c) a recent multimodal time series foundation model ChatTime [51].
For each time series model, we compare our TaTS framework against two baselines: (i) Numerical-only uni-modal modeling, which ignores the paired texts and utilizes only the numerical time series with the given time series model. (ii) MM-TSFLib  [35], a recently proposed library for multimodal time series forecasting that performs a linear interpolation between the output of the time series model and the text embeddings, treating the texts as a bag-of-words.

### v3

@@ Abstract
While many advances in time series models focus exclusively on numerical data, research on multimodal time series, particularly those involving contextual textual information, remains in its infancy. With recent progress in large language models and time series learning, we revisit the integration of paired texts with time series through the Platonic Representation Hypothesis (Huh et al., 2024), which posits that representations of different modalities converge to shared spaces. In this context, we identify that time-series-paired texts may naturally exhibit periodic properties that closely mirror those of the original time series. Building on this insight, we propose a novel framework, Texts as Time Series (TaTS), which considers the time-series-paired texts to be auxiliary variables of the time series. TaTS can be plugged into any existing numerical-only time series models and effectively enable them to handle time series data with paired texts. Through extensive experiments on both multimodal time series forecasting and imputation tasks across benchmark datasets with various existing time series models, we demonstrate that TaTS can enhance multimodal predictive performance without modifying model architectures. Our Code is available at https://github.com/iDEA-iSAIL-Lab-UIUC/TaTS.
@@ 1 Introduction
Figure 1: Mean Square Error of modeling frameworks of time series with paired texts. Full results in Appendix E.2.
Time series modeling plays an important role in a wide range of real-world applications, including finance (Sezer et al., 2020), healthcare (Zhang et al., 2024b), climate (Fu et al., 2024), and energy systems (Kotzur et al., 2017; Li et al., 2015). While extensive research has focused on approaches that rely solely on the numerical values of time series (Zhou et al., 2021; Liu et al., 2024c; Wu et al., 2023b; Wang et al., 2024c), real-world scenarios often involve additional modalities that co-occur with the time series and can provide valuable complementary information (De Baets and Harvey, 2023; Rai et al., 2023; Kyei and Antwi, 2017; Wang et al., 2024d).
In scenarios like pandemic policymaking, economic planning, or investment strategies, textual information can provide explanations, updates, or external factors that influence the underlying numerical patterns. However, research on effectively leveraging data from other modalities paired with time series remains in its early stages. In this work, we focus on time-series paired with texts at each timestamp, a common data format where textual descriptions are associated with time series at each timestamp in a parallel manner, as illustrated in Figure 3 (left). For instance, during a pandemic, infection rates are often accompanied by government announcements and news reports in real-time (Cinelli et al., 2020). On the one hand, numerical-only models overlook valuable contextual information that may influence or explain the patterns in the time series. On the other hand, the current state-of-the-art approach (Liu et al., 2024a) disregards the unique positional characteristics that time-series-paired texts may inherently possess. Such limitations raise a pivotal question:
@@
In this paper, motivated by the Platonic Representation Hypothesis (PRH) (Huh et al., 2024), we pioneer the exploration of effectively leveraging paired texts to enrich time series analysis. We identify an intriguing phenomenon, which we term “Chronological Textual Resonance” (CTR): depending on data quality, time-series-paired texts may exhibit periodic patterns that closely reflect the temporal dynamics of their corresponding numerical time series. More specifically, despite variations in expressions, the hidden representations of two periodicity-lagged texts associated with time series may demonstrate high similarities, revealing a deeper alignment between textual and numerical modalities. We attribute this phenomenon to the fact that the paired texts inherently evolve in response to the dynamics of the time series itself. Further, we introduce TT-Wasserstein, a new metric designed to measure the level of CTR and quantify alignment quality.
Building on these insights, we propose Texts as Time Series (TaTS), a simple yet effective framework for integrating paired texts to enhance multimodal time series modeling. Previous studies have shown that different variables in one multivariate time series exhibit similar periodicity properties (Zhang and Yan, 2023; Wang et al., 2024d; Yi et al., 2024), and CTR suggests that time-series-paired texts follow a similar pattern. This observation implies that paired texts can be considered as special auxiliary variables to augment the original time series. Motivated by this, TaTS first transforms the paired textual information into a lower-dimensional representation, then combines the original time series with the textual representations as new variables to form an augmented time series. This augmented time series is subsequently fed into existing time series models, allowing them to capture both numerical and textual dynamics. TaTS offers two key benefits: (i) it effectively captures the evolving positional characteristics of texts paired with a time series; and (ii) it functions as a plug-in module, maintaining compatibility with existing time series models. Empirically, the proposed TaTS achieves state-of-the-art performance on both forecasting and imputation tasks. Notably, we observe that a higher CTR level (i.e., a lower TT-Wasserstein) correlates with greater improvements compared to numerical-only modeling. In summary, our contributions are:
@@
  • Experiments on diverse benchmark datasets and multiple time series models demonstrate TaTS’s superior performance without requiring modifications to model architectures.
@@ 2 Preliminary
where $\mathbf{e}_{t}$ is the encoded text representation at time $t$, $d_{\text{text}}$ is text embedding dimension, and $\theta_{\text{text}}$ are encoder parameters. In this work, we leverage pre-trained large language models to encode the texts.
@@ 3 Chronological Textual Resonance (CTR)
The Platonic Representation Hypothesis. PRH (Huh et al., 2024) posits that different modalities describing the same object converge towards a shared, latent representation. Extending this hypothesis, if time series and paired text both describe the same changing event, their representations are dynamic projections from a common underlying source, and should exhibit similar periodic properties.
To illustrate this hypothesis for time series with paired texts, we analyze three real-world datasets (Liu et al., 2024a), including (i) Economy: The time series represents trade data of the U.S., while the texts describe the general economic conditions of the country. (ii) Social Good: The time series captures the unemployment rate in the U.S., and the texts include detailed unemployment reports. (iii) Traffic: The time series reflects monthly travel volume trends from the U.S. Department of Transportation, with corresponding texts derived from traffic volume reports issued by the same department.
For each dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, we employ Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) to analyze the frequency components of time series data and identify its dominant periodic components, illustrated by the blue curves in Figure 2. To examine the periodicity of texts, we embed each $s_{t}\in\mathbf{S}$ to obtain the text embedding $e_{t}$ at timestamp $t$. Then, we compute their lag-similarity, defined as $d_{l}=\sum_{t}\mathrm{cos}(e_{t},e_{t+L})$ where $L$ is the lag and $\mathrm{cos}(\cdot,\cdot)$ represents the cosine similarity. If the text embeddings exhibit a significant periodic pattern, the lag-similarity $d_{l}$ will also fluctuate periodically as the lag $l$ increases (proof in Proposition A.1). Finally, we identify major frequencies (with the largest amplitudes) of the texts by applying FFT to text lag-similarity, and mark them with red dashed lines, as shown in Figure 2. Detailed process is provided in Appendix A. We find that the major frequencies of the paired texts closely match those of the time series. Specifically, the paired texts also show periodicity of 12 (frequency 0.083) for monthly sampled time series, indicating that the paired texts exhibit periodicity that is strongly aligned with the temporal dynamics of the time series.
@@
Why may CTR happen? We present three key reasons for the observed alignment in periodicity between time series and their paired texts: (i) Shared External Drivers: Both the time series and their paired texts are often influenced by common external factors, such as seasonal changes, recurring events, or societal and economic cycles. These shared drivers naturally induce periodicity in both modalities. (ii) Influence of Time Series on Texts: Paired texts often serve as contextual reflections of the underlying time series, adapting and evolving in response to numerical trends. For instance, news articles or government reports accompanying economic indicators are frequently updated in response to the numerical trends. (iii) Texts Contain Additional Variables with Aligned Periodicity: Paired texts often include additional variables that are closely related to the time series. For example, if the time series represents economic GDP data, the accompanying texts may reference related variables such as stock market indices or inflation rates. These related variables often exhibit periodicity patterns aligned with the time series and affect the periodicity of the paired texts.
Table 1: TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets and their random shuffle.
@@
Quantifying CTR level. Of course we cannot expect every time series with paired texts at each timestamp to exhibit periodicity similarity. In some cases, texts aligned with timestamps may not provide meaningful or complementary information, for example, daily lottery winning numbers. To this end, we propose a new metric for CTR, TT-Wasserstein, defined as the Wasserstein distance between the normalized spectra of time series and texts. Formally, based on Wasserstein distance $W(P,Q)$ (Kantorovich, 1960), for time series with paired texts dataset $\mathcal{D}=\{\mathbf{X},\mathbf{S}\}$, after computing normalized frequencies and amplitudes $\tilde{f}_{\text{texts}},\tilde{f}_{\text{ts}},\tilde{a}_{\text{texts}},\tilde{a}_{\text{ts}}$,
Intuitively, $\text{TT-Wasserstein}(\mathcal{D})$ quantifies the discrepancy between the spectral distributions of paired texts and time series data. By design, a lower value of $\text{TT-Wasserstein}(\mathcal{D})$ should indicate a higher alignment between the textual and numerical modalities. To validate this, we compute our TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets, whose web-retrieved texts are filtered, disentangled, and summarized for relevance and alignment. We also compute TT-Wasserstein on time-series-shuffled or text-shuffled versions of the datasets to disrupt alignment. As shown in Table 1, the corrupted Time-MMD datasets yield much larger Wasserstein distances compared to the original datasets. This finding suggests that TT-Wasserstein can serve as an indicator for the alignment between modalities and a gauge of dataset quality. In Section 5, we will show that TT-Wasserstein can even predict the potential effectiveness of our proposed framework. Additionally, because TT-Wasserstein is an empirical statistical metric, we further analyze its sensitivity and estimation stability in Appendix A.1.
@@ 4 Texts as Time Series
An overview of the proposed Texts as Time Series (TaTS) is illustrated in Figure 3.
Concurrent Texts are Secretly Auxiliary Variables. As we elucidated in Section 3, the properties of concurrent texts closely align with those of variables in a multivariate time series: similar to numerical variables, concurrent texts are influenced by shared external drivers and interact dynamically with the time series. Furthermore, mapping concurrent texts to structured variables enables capturing hidden variables embedded within the concurrent texts.
Given the dataset $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}\}$, TaTS first embed the texts using a text encoder $\mathcal{H}_{\text{text}}$ to obtain text embeddings $\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\in\mathbb{R}^{d_{\text{text}\times T}}$. Since the text embedding dimension $d_{\text{text}}$ is typically much larger than the number of variables in the time series, we reduce the dimensionality of the text embeddings by applying a Multi-Layer Perceptron (MLP), mapping them into a lower-dimensional space of reduced dimensionality $d_{\text{mapped}}$.
$\mathbf{z}_{t}=\text{MLP}(\mathbf{e}_{t};\theta_{\text{MLP}})\in\mathbb{R}^{d_{\text{mapped}}}$ (5)
@@
$\mathbf{U}=[\mathbf{X};\mathbf{Z}^{\intercal}]_{\text{dim=1}}\in\mathbb{R}^{T\times(N+d_{\text{mapped}})}$ (6)
@@
$\widehat{\mathbf{X}}_{T+1:T+H}=\mathcal{F}\left(\mathbf{U}_{1:T};\theta_{\text{forecast}}\right)[:N]\in\mathbb{R}^{H\times N}$ (7)
where $\mathcal{F}(\cdot;\theta_{\text{forecast}})$ denotes the time series forecasting model with parameters $\theta_{\text{forecast}}$, and $[:N]$ extracts the first $N$ variables corresponding to the original time series.
@@
$\mathcal{L}_{\text{forecast}}(\mathbf{X},\widehat{\mathbf{X}})=\frac{1}{H\cdot N}\sum_{t=T+1}^{T+H}\sum_{i=1}^{N}\left(\mathbf{X}_{t,i}-\widehat{\mathbf{X}}_{t,i}\right)^{2}$ (8)
@@ 5 Experiment
Table 2: Time-series forecasting performance. Compared to numerical-only unimodal modeling and $M$-TSFLib, TaTS consistently and significantly enhances existing time series models to effectively handle paired texts. The results are averaged across all prediction lengths. Full results reported in Appendix E.3. Promotion (positive or negative) denotes the percentage reduction (or increase) in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
[…]
In this section, we empirically validate the effectiveness of the proposed TaTS framework. In particular, we use GPT2 (Radford et al., 2019) encoder to embed the paired texts. We also validate the performance of TaTS with other text encoders across different datasets in section 5.2 and Appendix E.7. Implementation details are provided in Appendix D.4.
Datasets. We evaluate our framework on 18 real-world datasets from Time-MMD (Liu et al., 2024a), FNSPID Dong et al. (2024), and FNF Wang et al. (2024b). The datasets span diverse domains with sample frequencies ranging from daily to weekly and monthly. More details are in Appendix D.1.
Time Series Models and Baselines. To demonstrate the compatibility of TaTS with existing time series models, we integrate TaTS with 9 widely used models across different categories, including (i) Transformer-based models: iTransformer (Liu et al., 2024c), PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021) and Transformer (Vaswani et al., 2017). (ii) Linear models: DLinear (Zeng et al., 2023). (iii) Frequency-based models: FEDformer (Zhou et al., 2022b), FiLM (Zhou et al., 2022a).
We compare our TaTS framework with a wide array of baselines, including (a) numerical-only uni-modal modeling, which ignores the paired texts and utilizes only the numerical time series with the given time series model; (b) MM-TSFLib  (Liu et al., 2024a), a recently library proposed together with Time-MMD datasets for multimodal time series forecasting; (c) covariate-based methods: N-BEATS (Oreshkin et al., 2020), N-HiTS (Challu et al., 2022); (d) convolution-based method TCN (Bai et al., 2018); (e) a recent multimodal time series foundation model ChatTime (Wang et al., 2025).